OpenStudy (anonymous):

The sum of two numbers is x. If one of the numbers is 12, what is the other? 12 -x x- 12 12 +x 12x

OpenStudy (anonymous):

a + b = x if a is 12 then: 12 + b = x b = x - 12

OpenStudy (anonymous):

thank you for that